a man pays $8 for a hot dog and a soda. The hot dog is $2 more than the soda. How much does each cost?
Answer by stanbon(57967)

a man pays $8 for a hot dog and a soda.
The hot dog is $2 more than the soda.
How much does each cost?
---------
Equations:
h + s = 8 dollars
h = s+2
-----
Substitute for "h" and solve for "s":
s+2 + s = 8
2s = 6
s = 3 dollars (price of a soda)
---
Solve for "h":
h + s = 8
h + 3 = 8
h = 5 dollars (price of a hot dog)
======================================
Cheers,
Stan H.